# Victoria, BC, Canada

### v3

@@ 5.2 Sam’s Study: Studying The Impact of a Generative AI Tool that Automates Coding Reviewing
Sam (they, them) is an early stage Ph.D. student collaborating with a large company that is trialing a new Generative AI tool to automate part of the code review process. Sam’s initial research goal is to test their hypothesis that a new GenAI tool that automates the code review process will improve the time to merge of code changes and improve the quality of the reviewed code (by reducing the number of bugs in the newly committed code).
@@
From the interviews combined with the mined data, novel insights about the tool’s efficacy and how it was used, emerged. Notably, from the sequential integration of the mined data and interviews, Sam learns that the developers are able to work more efficiently but they are potentially more complacent when using the AI reviewing tool. Sam reflects on the mixed insights and their implications, limitations, and future directions. For instance, these integrated insights can be used to shape code review tool design accordingly, by adding features to prompt developers to avoid complacency.
@@
A mixed methods design was not intended up front, but was an emergent or organic research design decision to explain the surprising results from the data mining. In addition to ensure the mining process and interviews were conducted with rigor, Sam must ensure that the sequential integration step was also done in a way to reliably explain the surprising result from the mined data. As Sam may be biased following this unexpected finding, they recruit another student to conduct these interviews. Note Sam, could have used an embedded design with a minor survey strand to explain the insights from the mined data but the findings were so unexpected they anticipated that interviews could also inform how the review tool could be improved before more mining.
@@
5.3 Vicki’s Study: Understanding the Adoption of a New Onboarding Process in a Large Scale Open Source Project
@@
The choice of a convergent parallel design supports Vicki to inductively analyze the communication about pairing qualitatively and link the communication instances to the commits that are potentially happening in parallel (or soon after). This design allows her to integrate and build a story about how the mentors and new contributors communicate and pair, and the impact of that communication on the timing and presence of subsequent or in parallel commits. She considers the quantitative aspects (number and timing) of the commit data in an inductive way to see what emerges (she has no initial hypothesis about the impact of this informal pairing over Slack). As she integrates the two data sets, she is careful to specify the timing intervals as she links the two sets of data (the timing interval used is suggested by other research that links these two types of data). Together, these two data sources lead to complementary insights and theoretical insights about the nature of the communication about pairing and the commits the communication is connected to. Her analysis leads to meta-inferences about the pairing activities and commit behaviour, notably that pairing during onboarding lead to faster commits. The rigor of her design is increased by checking the coding reliability with other researchers and running statistical analysis on the quantitative data she collects.
@@
Vicki seeks permission from her research participants and project owners to both monitor the project communication channels and mine logged data about code commits and their comments. She also ensures that new contributors agree to having their communications with the mentors analyzed for the study, and that they may opt out at any time (ensuring any collected information is deleted from the study data across both methods). Likewise, mentors also must give permission and can opt out at any time. She anonymizes the identity of the participants she observes in the communication channel, ensuring no identifying data remains in the reported data. Additionally, during the results presentation, she aggregates the participant’s profile to avoid putting any particular group in evidence, as this was not part of her study’s goal (e.g., gathering only relevant personal data for her research and not showing personal characteristics in her participant’s demographic table). Vicki recognizes that mining the commit data comes with the potential risk of introducing significant privacy concerns, so she also obfuscates the commit data. Vicki conducts her observations of the Slack channels in a transparent but nonobtrusive manner, ensuring security in its storage and planned deletion after the study.
@@ 5.4 Zara’s Study: Evaluating the Effectiveness of an Automated tool for Generating Tests
Zara’s study reveals that using the automated tool for test generation not only improves the speed of test creation compared to standard IDE features (identified from the experiment) but through the embedded survey she finds it also enhances the participants’ confidence in the tests they generate and their overall understanding of software testing concepts. Using a data-based integration that involes the simultaneous analysis of both quantitatve and qualitate data brings this new meta-inference about the impact of this new tool (when compared to the features in an IDE) on student perceived confidence and learning.
@@
The use of student participants calls for several ethical considerations. Participation is voluntary and their identity is kept confidential across both methods. Zara takes care during recruitment that the students’ teachers are not involved in the study or research, that the teachers do not know who participates in the study, and that they do not put any real or apparent pressure on students to participate and thus don’t introduce power dynamics. The students are reassured that their decision to participate (or not) and nature of the feedback they provide will have no bearing on their performance in the class. The students’ time to participate is also valued. The potential benefit of gaining experience with test generation is communicated to the students during recruitment (this value was validated with teachers before the study). Care is taken to ensure students do not feel stressed during the study, and they can opt out at any time.
@@ 5.5 Learning from Other Examples
So far, we have shared that using MMR enables SE researchers to study their research topics from a more holistic perspective, expanding and complimenting methods to achieve more and novel insights. We also note from the scenarios we provide (and the paper examples we reference in Section 4.3) that mixed methods are particularly valuable in studying socio-technical aspects of SE phenomena. The research scenarios we provided help to demonstrate the broad and varied landscape of possible MMR designs in SE research. The ACM SigSoft Empirical Standards for Multiple and Mixed Methods also provides a list of example MMR papers sigsoft , but care has to be taken in using these and other examples of MMR designs as “exemplars” to guide future work. With this in mind, we next present “antipatterns” to avoid or to watch out for when designing or critiquing an MMR design. An earlier version of these antipatterns also appears in the ACM SigSoft Standards, contributed by the first author of this paper.
@@ 6 Antipatterns of Mixed Method Research Designs
The principles and research design patterns we described above can be followed when designing a study. And through the scenarios above, we illustrated how the principles of MMR could be followed. In the following, we provide examples of “what not to do” and “what to avoid”.
As researchers, authors and reviewers of papers, we have noted a number of common antipatterns in how mixed methods are used or presented. Some of the antipatterns we encountered originate in issues with how the research is presented, rather than inherent to the conduct of the research. We introduce presentation antipatterns first noting how they can be potentially fixed. The other study design antipatterns are harder to “fix” after the research has been conducted, but hopefully can be identified at design time (or during review of a paper). We recommend researchers think about these antipatterns early in their research process.
@@ 6.1 Presentation Antipatterns
As mentioned, these antipatterns are concerns in how the research is presented or written. They are for the most part relatively easy to fix.
  • Uninvited guest or party crasher: This antipattern occurs when one of the methods used is not clearly introduced in a paper introduction or methodology and makes an unexpected entrance in the later part of the paper, such as in the discussion or limitation sections. To fix this, the author should clearly mention all methods in the introduction and methodology sections of the paper, justifying why they were used. Even with emergent designs, as in the case of Sam (Sec. 5.2), all methods used should be presented early in the paper with a mention that the design was emergent.
  • Smoke and mirrors: This antipattern can be spotted when a paper attempts to oversell a study as an MMR when one of the methods merely offers a token or anecdotal contribution to the research motivation or findings. This can be fixed by clarifying the presentation in the paper. For example, in the case of the embedded design carried out by Zara (Sec. 5.4), if the treatment of the embedded part of the study (to find out how the tool impacted confidence in testing) was shallow and the number of participants surveyed was very low, the paper would have oversold the contribution as a MMR study.
@@ 6.2 Study Design Antipatterns
These antipatterns are ingrained in how the study was designed or conducted. They are hard if not impossible to fix afterwards but thinking about them early on can help researchers avoid them.
@@
  • Missing the mark: Misalignment of MMR design with the research question or objective. This can be seen when one (or even both) of the methods do not make sense in light of the research question being posed. The fix for this should happen at the study design phase by asking if every method answers the core research questions being asked. In the case of Zara, the research question considered not just the testing speed, but also testing confidence. If the embedded method did not lead to insights about testing confidence through the data integration step, this design could have been described as “missing the mark”.
@@
  • Selling your soul: This antipattern is related to “Missing the mark” and may occur when a researcher employs an additional method to appeal to a future reviewer who they believe has preference for a particular methodology or study outcome, but the use of the method does not contribute substantively to the research findings. This can be addressed by using the design one authentically believes in, or by ensuring that the researcher draws out meta-inferences from the different methods used. In the case of Ali (Sec. 5.1), the findings from the interviews may have been sufficient to stand on their own merit. But Ali may also have been influenced to do a quaNsurvey to appease to a reviewer that may be focused on knowing if the findings generalized to more teams in the company.
@@
  • Cargo cult research [^11] Cargo Cult can be a metaphor for superficial imitation of a process without basic understanding of its mechanism.: Using research methods in which the research team lacks expertise but that they have observed being used to generate “exciting findings” in well cited papers. Collaborating with other researchers with broader expertise may help avoid this antipattern. For example, in the case of Vicki adding a data mining method to the observation part of the design may require working with collaborators with mining expertise if Vicki lacks that expertise.
@@
  • Sample contamination: This antipattern may occur in an MMR sequential design where the same participants are used in multiple, sequential methods without accounting for potential contamination from earlier method(s) to later ones (where participants have already seen an intervention or know about the research). For example, in the case of Ali, Ali should be careful not to survey the participants already interviewed. However, in some cases the design itself may involve selecting participants (e.g., for interviews) from an earlier method (e.g., a survey) and the resulting “contamination” is intentional (for example, in the case of a embedded design). Any possible contamination should be justified or reported, and mentioned in limitations.
@@
  • Lost opportunity: This may occur in an MMR sequential design, when there is a failure to use findings from an earlier method during the development of an instrument for a later method. This should be caught during the implementation of the study as it cannot be fixed through the presentation (alternatively, the studies can be written up as a two-phase study or as two separate studies with two different contributions). In the case of Ali, not using insights from the interviews to ask pertinent questions in the survey about the security tool mandate could be a lost opportunity.
@@
  • Questionable ethics: Finally, even when all of the above antipatterns are avoided, i.e., the MMR study is applied and presented well, it may come across as having questionable aims, or as using questionable means to achieve the aims. The researcher should clearly state the motivation for the study and expected benefit to society or research to help justify the possible risks a study may involve. In the cases of all the fictional researchers above, we mentioned at least some of the ethical issues to be considered. All too often though we see papers were the respondents are not treated with this same care. We refer the reader to the points raised about conducting MMR ethically in Section 3.4 and encourage all studies should undergo an ethical review.
@@ 7 Discussion
In this section, we address a number of issues and nuances related to MMR. We discuss MMR synergies with other methods, including difference between mixed and multi methods research. Finally, we present a call for action.
@@ 7.1 Synergies with Other Methods
Mixed methods research can easily be confused with multi methods research [^12] We remind the reader this and other terms are defined in the Glossary found in the Appendix. The difference between mixed methods and multi methods lies in what types of data are used and how the data is used. Mixed methods requires the integration of quantitative (quaN) and qualitative (quaL) data. Multi methods can be applied with any combination of quaN-quaN or quaL-quaL methods. The principles we outline in Section 3 all apply to multi method research as well but we focus this paper on the use of the designs for quaN and quaL methods which can bring more emphasis on the socio-technical aspects in SE research.
@@ 7.2 A Call to Action
Much of SE research is inherently socio-technical and as such can benefit from the use of Mixed Methods to bring insights on the underlying context and to strengthen the theoretical contributions from our work, and capability to generalize or translate our findings to other contexts. We call for a principled use of mixed methods in SE research. Figure 7 summarizes the key points we raised in our paper: the MMR definition, a landscape of MMR designs, four principles to guide MMR in SE, and a list of antipatterns of MMR designs to be considered.
@@
In adapting the guiding principles we propose for an SE context, we do not simply promote MMR but emphasize using MMR ethically, as captured in the principle of ethical research. As a research community we have conducted large-scale surveys with developers and mined developer forums which in some cases has led to outrage in the developer community and introduction of new policies for unsolicited research calls for participation (see Section 3.4). As researchers, we should have empathy for the people we are studying, and indeed aim to help. They are real people. This is another reason why we include examples of fictional researchers, painting a picture of these researchers as human beings and the struggles they may face. Many of these struggles are not shared in the articles they write but perhaps could be so we can learn from each other. We recommend using MMR to make SE research more human focused–amplifying the human side of the quantitative data with the inclusion of qualitative data, both collected ethically.
As the lure of AI tools tempts the SE research community with promises of easing various research tasks and methods, it is timely that we get into the habit of asking ourselves ‘why’, and not just be driven by the temptation of ‘why not’, as captured in our principle of methodological rationale. At the same time, as we begin to explore the risk and the opportunities of using LLMs for qualitative research bano2023LLM ; hoda2024Book , let us remind ourselves of the value of ‘deep work’, captured in the principle of procedural rigour. We strongly believe that using MMR studies conducted in interdisciplinary research settings can help address complex, socio-technical research questions and drive meaningful research, captured in our principle of novel integrated insights.
These guiding principles for using MMR in SE are required now more than ever. We hope this first step will help present and future SE researchers use mixed methods research ethically, rationally, and rigorously to generate novel insights, and encourage our community to develop more prescriptive guidelines and playbooks for using and evaluating mixed methods research in SE.

### v4

@@ 5.2 Sam’s Study: Studying The Impact of a Generative AI Tool that Automates Coding Reviewing
Sam (they, them) is an early-stage Ph.D. student collaborating with a large company that is trialing a new Generative AI tool to automate part of the code review process. Sam’s initial research goal is to test their hypothesis that a new Generative AI (GenAI) tool that automates the code review process will improve the time to merge of code changes and improve the quality of the reviewed code (by reducing the number of bugs in the newly committed code).
@@
From the interviews combined with the mined data, novel insights about the tool’s efficacy and how it was used, emerged. Notably, from the sequential integration of the mined data and interviews, Sam learns that the developers are able to work more efficiently but they are potentially more complacent when using the GenAI reviewing tool. Sam reflects on the mixed insights and their implications, limitations, and future directions. For instance, these integrated insights can be used to shape code review tool design accordingly, by adding features to prompt developers to avoid complacency.
@@
A mixed methods design was not intended up front, but was an emergent or organic research design decision to explain the surprising results from the data mining. In addition to ensure the mining process and interviews were conducted with rigor, Sam must ensure that the sequential integration step was also done in a way to reliably explain the surprising result from the mined data. As Sam may be biased following this unexpected finding, they recruit another student to conduct these interviews. Note Sam could have used an embedded design with a minor survey strand to explain the insights from the mined data but the findings were so unexpected they anticipated that interviews could also inform how the review tool could be improved before more mining.
@@
5.3 Vicki’s Study: Understanding the Adoption of a New Onboarding Process in a Large-Scale Open Source Project
@@
This design allows Vicki to integrate and build a story about how the mentors and new contributors communicate and pair, and the impact of that communication on the timing and presence of subsequent or in parallel commits. She considers the quantitative aspects (number and timing) of the commit data in an inductive way to see what emerges (she has no initial hypothesis about the impact of this informal pairing over Slack). As she integrates the two data sets, she is careful to specify the timing intervals as she links the two sets of data (the timing interval used is suggested by other research that links these two types of data). Together, these two data sources lead to complementary insights and theoretical insights about the nature of the communication about pairing and the commits the communication is connected to. Her analysis leads to meta-inferences about the pairing activities and commit behaviour, notably that pairing during onboarding lead to faster commits. The rigor of her design is increased by checking the coding reliability with other researchers and running statistical analysis on the quantitative data she collects.
@@
Vicki seeks permission from her research participants and project owners to both monitor the project communication channels and mine logged data about code commits and their comments. She also ensures that new contributors agree to having their communications with the mentors analyzed for the study, and that they may opt out at any time (ensuring any collected information is deleted from the study data across both methods). Likewise, mentors also must give permission and can opt out at any time. She anonymizes the identity of the participants she observes in the communication channel, ensuring no identifying data remains in the reported data. Vicki recognizes that mining the commit data comes with the potential risk of introducing significant privacy concerns, so she also obfuscates the commit data. Vicki conducts her observations of the Slack channels in a transparent but nonobtrusive manner, ensuring security in its storage and planned deletion after the study.
@@ 5.4 Zara’s Study: Evaluating the Effectiveness of an Automated tool for Generating Tests
Zara’s study reveals that using the automated tool for test generation not only improves the speed of test creation compared to standard IDE features (identified from the experiment) but through the embedded survey she finds it also enhances the participants’ confidence in the tests they generate and their overall understanding of software testing concepts. Using a data-based integration that involves the simultaneous analysis of both quantitative and qualitative data brings this new meta-inference about the impact of this new tool (when compared to the features in an IDE) on student perceived confidence and learning.
@@
The use of student participants calls for several ethical considerations. Participation is voluntary and their identity is kept confidential across both methods. Zara takes care during recruitment that the students’ teachers are not involved in the study or research, that the teachers do not know who participates in the study, and that they do not put any real or apparent pressure on students to participate and thus do not introduce power dynamics. The students are reassured that their decision to participate (or not) and nature of the feedback they provide will have no bearing on their performance in the class. The students’ time to participate is also valued. The potential benefit of gaining experience with test generation is communicated to the students during recruitment (this value was validated with teachers before the study). Care is taken to ensure students do not feel stressed during the study, and they can opt out at any time.
@@ 5.5 Learning from Other Examples
So far, we have shared that using MMR enables SE researchers to study their research topics from a more holistic perspective, expanding and complementing methods to achieve more novel insights. We also note from the scenarios we provide (and the paper examples we reference in Section 4.3) that mixed methods are particularly valuable in studying socio-technical aspects of SE phenomena. The research scenarios we provided help to demonstrate the broad and varied landscape of possible MMR designs in SE research. The ACM SIGSOFT Empirical Standards for Multiple and Mixed Methods also provides a list of MMR papers ralph2020empirical , but care has to be taken in using these and other examples of MMR designs as “exemplars” to guide future work. With this in mind, we next present “antipatterns” to avoid or to watch out for when designing or critiquing an MMR design. An earlier version of these antipatterns also appears in the ACM SIGSOFT Empirical Standards, contributed by the first author of this paper.
@@ 6 Antipatterns of Mixed Method Research Designs
The principles and research design patterns we described above can be followed when designing a study. And through the scenarios, we illustrated how the principles of MMR could be followed. In the following, we provide examples of “what not to do” and “what to avoid”.
As researchers, authors and reviewers of papers, we have noted a number of common antipatterns in how mixed methods are used or presented. Some of the antipatterns we encountered originate in issues with how the research is presented, rather than inherent to the conduct of the research. We introduce presentation antipatterns first noting how they can be potentially fixed. The other study design antipatterns are harder to “fix” after the research has been conducted, but hopefully can be identified at design time (or during review of a paper). We recommend researchers consider these antipatterns early in their research process.
@@ 6.1 Presentation Antipatterns
As mentioned, these antipatterns are concerns in how the research is presented or written.
  • Uninvited guest or party crasher: This antipattern occurs when one of the methods used is not clearly introduced in a study report introduction or methodology and makes an unexpected entrance in the later part of the report, such as in the discussion or limitation sections. To fix this, the author should clearly mention all methods in the introduction and methodology sections of the study report, justifying why they were used. Even with emergent designs, as in the case of Sam (Sec. 5.2), all methods used should be presented early in the paper with a mention that the design was emergent.
  • Smoke and mirrors: This antipattern can be spotted when a paper attempts to oversell a study as an MMR when one of the methods merely offers a token or anecdotal contribution to the research motivation or findings. This can be fixed by clarifying the presentation in the paper. For example, in the case of the embedded design carried out by Zara (Sec. 5.4), if the treatment of the embedded part of the study (to find out how the tool impacted confidence in testing) was shallow and the number of participants surveyed was very low, the paper would have oversold the contribution as an MMR study.
@@ 6.2 Study Design Antipatterns
These antipatterns are ingrained in how the study was designed or conducted. They are hard, if not impossible to fix afterwards, but thinking about them early on can help researchers avoid them.
@@
  • Missing the mark: This antipattern occurs when the MMR design is not aligned with the research question or objective. This can be seen when one or more of the methods do not make sense in light of the research question being posed. The fix for this should happen at the study design phase by asking if every method answers the core research questions being asked. In the case of Zara, the research question considered not just the testing speed, but also testing confidence. If the embedded method did not lead to insights about testing confidence through the data integration step, this design could have been described as “missing the mark”.
@@
  • Selling your soul: This antipattern is related to “missing the mark” and may occur when a researcher employs an additional method to appeal to a future reviewer who they believe has preference for a particular methodology or study outcome, but the use of the method does not contribute substantively to the research findings. This can be addressed by using the design one authentically believes in, or by ensuring that the researcher draws out meta-inferences from the different methods used. In the case of Ali (Sec. 5.1), the findings from the interviews may have been sufficient to stand on their own merit. But Ali may also have been influenced to do a quaN survey only to appeal to a reviewer that may be focused on knowing if the findings generalized to more teams in the company.
@@
  • Cargo cult research [^11] Cargo cult can be a metaphor for superficial imitation of a process without basic understanding of its mechanism.: Using research methods in which the research team lacks expertise but that they have observed being used to generate “exciting findings” in well cited papers that use these methods. Collaborating with other researchers with broader expertise may help avoid this antipattern. For example, in the case of Vicki, adding a data mining method to the observation part of the design may require working with collaborators with mining expertise if Vicki lacks that expertise.
@@
  • Sample contamination: This antipattern may occur in an MMR sequential design where the same participants are used in multiple, sequential methods without accounting for potential contamination from earlier method(s) to later ones (where participants have already seen an intervention or know about the research). For example, in the case of Ali, Ali should be careful not to survey the participants already interviewed. However, in some cases the design itself may involve selecting participants (e.g., for interviews) from an earlier method (e.g., a survey) and the resulting “contamination” is intentional (for example, in the case of an embedded design). Any possible contamination should be justified or reported, and mentioned in the limitations of the study report.
@@
  • Lost opportunity: This may occur in an MMR sequential design when there is a failure to use findings from an earlier method during the development of an instrument for a later method. This should be caught during the implementation of the study as it cannot be fixed through the presentation (alternatively, the studies can be written up as a two-phase study or as two separate studies with two different contributions). In the case of Ali, not using insights from the interviews to ask pertinent questions in the survey about the security tool mandate could be a lost opportunity.
@@
  • Questionable ethics: Finally, even when all of the above antipatterns are avoided, i.e., the MMR study is applied and presented well, it may come across as having questionable aims, or as using questionable means to achieve the aims. The researcher should clearly state the motivation for the study and expected benefit to society or research to help justify the possible risks a study may involve. In the cases of all the fictional researchers above, we mentioned at least some of the ethical issues to be considered. We refer the reader to the points raised about conducting MMR ethically in Section 3.4 and encourage all studies should undergo an ethical review.
@@ 7 Discussion
In this section, we discuss how MMR relates to other methods (such as multi-methods) and present a call for action.
@@ 7.1 Synergies with Other Methods
Mixed methods research can easily be confused with multi-methods research. [^12] We remind the reader this and other terms are defined in the Glossary found in the Appendix The difference between mixed methods and multi-methods lies in what types of data are used and how the data is used. Mixed methods requires the integration of quantitative (quaN) and qualitative (quaL) data. Multi-methods can be applied with any combination of quaN-quaN or quaL-quaL methods. The principles we outline in Section 3 all apply to multi-method research as well.
@@ 7.2 A Call to Action
Much of SE research is inherently socio-technical and as such can benefit from the use of MMR to bring insights on the underlying context and to strengthen the theoretical contributions from our work, and capability to generalize or translate our findings to other contexts. We call for a principled use of mixed methods in SE research. Figure 7 summarizes the key points we raised in our paper: the MMR definition, a landscape of MMR designs, four principles to guide MMR in SE, and a list of antipatterns of MMR designs to be considered.
@@
In adapting the guiding principles we propose for an SE context, we do not simply promote MMR but emphasize using MMR ethically, as captured in the principle of ethical research. As a research community we have conducted large-scale surveys with developers and mined developer forums which in some cases has led to outrage in the developer community and introduction of new policies for unsolicited research calls for participation (see Section 3.4). As researchers, we should have empathy for the people we are studying, and indeed aim to help them. They are real people. This is another reason why we include examples of fictional researchers, painting a picture of these researchers as human beings and the struggles they may face. Many of these struggles are not shared in the articles they write but perhaps could be so we can learn from each other. We recommend using MMR to make SE research more human focused–amplifying the human side of the quantitative data with the inclusion of qualitative data, both collected ethically.
As the lure of AI tools tempts the SE research community with promises of easing or accelerating various research tasks and methods, it is timely that we get into the habit of asking ourselves ‘why’, and not just be driven by the temptation of ‘why not’, as captured in our principle of methodological rationale. At the same time, as we begin to explore the risk and the opportunities of using LLMs for qualitative research bano2023LLM ; hoda2024Book , let us remind ourselves of the value of ‘deep work’, captured in the principle of procedural rigor. We strongly believe that using MMR studies conducted in interdisciplinary research settings can help address complex, socio-technical research questions and drive meaningful research, captured in our principle of novel integrated insights.
These guiding principles for using MMR in SE are required now more than ever. We hope this first step will help present and future SE researchers to use mixed methods research ethically, rationally, and rigorously to generate novel insights, and encourage our community to develop more prescriptive guidelines and playbooks for using and evaluating mixed methods research in SE.